OpenStudy (anonymous):

Solve the proportion using cross products. 0.5 = 0.03 --- --- 5 n

zepdrix (zepdrix):

Giving us,\[\large 0.5 \times n=0.03 \times 5\]

zepdrix (zepdrix):

So the right side will multiply to give us,\[\large 0.5n=.15\]We'll divide both sides by 0.5,\[\large \frac{\cancel{0.5}n}{\cancel{0.5}}=\frac{.15}{0.5}\]On the left, the 0.5's will cancel out, giving us 1.

zepdrix (zepdrix):

So we get a final answer of ... whatever this gives you c: Throw it into your calculator. Lemme know if yer way confused about any of that. It's a tricky first step to get used to. \[\large n=\frac{.15}{.5}\]

OpenStudy (anonymous):

So I'd come up with 0.3 for the answer? thanks!

zepdrix (zepdrix):

Yup sounds right \c:/